OpenStudy (anonymous):

Find the slant height of a cone with a surface area of 282.74 ft. sq. and a diameter of 10 ft. A)13 B)14 C)12 D)15

OpenStudy (anonymous):

Surface Area Of Cone = \[\pi rl + \pi r^2\] where, l is Slant height and r is radius.. D= 10ft, So r = 10/2 = 5.. SA = 282.74 So, 282.74 = (3.14)(5)*l + (3.14)(5)(5) 282.74 - 78.5 = 15.7*l l = 204.24/15.7 = 13.00 (approx).. So, the Slant Height is 13.. Answer A)...

OpenStudy (anonymous):

thats alot of work ! thank you again.

OpenStudy (anonymous):

Welcome dear...